# Holographic Integrated Data and Energy Transfer

### v1

@@ Abstract
Thanks to the application of metamaterials, holographic multiple-input multiple-output (H-MIMO) is expected to achieve a higher spatial diversity gain by enabling the ability to generate any current distribution on the surface. With the aid of electromagnetic (EM) manipulation capability of H-MIMO, integrated data and energy transfer (IDET) system can fully exploits the EM channel to realize energy focusing and eliminate inter-user interference, which yields the concept of holographic IDET (H-IDET). In this paper, we invetigate the beamforming designs for H-IDET systems, where the sum-rate of data users (DUs) are maximized by guaranteeing the energy harvesting requirements of energy users (EUs). In order to solve the non-convex functional programming, a block coordinate descent (BCD) based scheme is proposed, wherein the Fourier transform and the equivalence between the signal-to-interference-plus-noise ratio (SINR) and the mean-square error (MSE) are also conceived, followed by the successive convex approximation (SCA) and an initialization scheme to enhance robustness. Numerical results illustrate that our proposed H-IDET scheme outperforms benchmark schemes, especially the one adopting traditional discrete antennas. Besides, the near-field focusing using EM channel model achieves better performance compared to that using the traditional channel model, especially for WPT where the EUs are usually close to the transmitter.
@@ I-A Background
In the era of next Internet of Things (IoT), ubiquitous intelligent connections are establishing between people and things. Deploying millions or even billions of IoT devices having low hardware size, complexity and power consumption into smart cities will readily breeds some emerging applications such as intelligent transportation, virtual reality and cloud computing [1]. However, it is posing a challenge for the energy requirement of all the IoT devices, where manually replacing or charging batteries requires much high maintenance costs and even brings about some safety hazards. In the Ambient IoT, the low-power devicies are expected to harvest energy by light, motion, heat and especially radio frequency (RF) signals [2]. By utilizing existing transmitters such as cellular base stations, TV towers, and Wi-Fi access points, RF based wireless energy transfer (WET) has the advantages of better controllability and stability [3]. By coordinating with wireless data transfer (WDT), integrated data and energy transfer (IDET) is a promising technology in the future to provide both energy self-sustaining and data communication services towards low-power devices in IoT [4].
@@
Moreover, thanks to the ability to manipulate the EM field, H-MIMO can not only ultimately approaching the communication limit but also achieve unprecedented concentration of EM energy, which is benefitial to both the WDT and WET [7]. With the aid of the propagation characteristics of the EM channel, the unprecedented energy concentration and multiplexing are realized for EUs, while the inter-user interference is well eliminated for DUs, which yields the concept of Holographic IDET (H-IDET).
@@ I-B Related works
In the multi-user IDET systems, the interference signals are usually harmful for WDT but can improve the WET performance. Therefore, the transceivers should be carefully designed to facilitate efficient IDET. Specifically, Xiong et. al. [8] investigated a broadcasting IDET system and analyzed the tradeoff between energy transfer versus information rate under nonlinear energy harvesting model. Yue et. al. [9] studied a millimeter wave IDET multicast system, where the data rate was maximized by satisfying wireless charging requirement with the aid of hybrid transceiver design. Large-scale antenna array are used in IDET to alleviate the inherent power attenuation caused by path loss in the channel. Gao et. al. [10] investigated an active intelligent reflecting surface (IRS) aided IDET system, where the beamforming designs were obtained by iterative algorithms, in oder to maximize the sum-rate of DUs and sum-power harvested by EUs. Zhang et. al. [11] considered an IRS assisted IDET system, where a deep reinforcement learning-based approach is proposed by utilizing proximal policy optimization framework, in order to maximize the energy efficiency.
@@
Moreover, there are lots of works studying H-MIMO. The metamaterial element of H-MIMO should satisfy the frequency response of Lorentzian form, which leads to different hardware structures [12]. Precisely, Deng et. al. [13] proposed a hardware design for H-MIMO to achieve amplitude-controlled beamforming, where a prototype was implemented to support real-time video transmission. Based on the amplitude-only-controlled hardware structure, they investigated the H-MIMO for cell communication systems [14], low-earth-orbit satellite systems [15] and integrate sensing and communication (ISAC) systems [5]. Besides, different from amplitude-control, the elements of the phase-only-controlled H-MIMO has Lorentzian-constrained phase model while keeing the amplitude constant, which can achieve considerable performance [16, 17].
@@
Furthermore, in order to explore the performance limits enabled by H-MIMO, EM theories should be applied to investigate the technique including physical antenna design, communication capacity analysis and channel modeling. Specifically, Badawe et. al. [18] proposed a practical physical architecture for the H-MIMO, where experimental results of the fabricated prototype matches well with numerical simulations. Besides, Hu et. al. [19] cosidered a communication system with a H-MIMO transmitter and several single-antenna receivers, where the normalized capacity and Degree of Freedom (DoF) measured per unit surface or per unit volume was analyzed. Moreover, Dardari et. al. [20] investigated the communication between two H-MIMO transceivers and the spatial DoF is derived in analytical expression. Furthermore, Pizzo et. al. [21] investigated the channel model of H-MIMO in scattering environments, where the model is further applied to evaluate the capacity of H-MIMO communication. In recent years, the EM based H-MIMO transceiver designs for communication systems are up to date. Specifically, Sanguinetti et. al. [22] investigated the line-of-sight communication betweem two H-MIMO transceivers in the nearfield. By utilizing Fourier basis functions, a multiplexing scheme was designed to generate orthogonal beamformers. Besides, Zhang et. al. [23] considered a multi-user H-MIMO system, where a pattern-division multiplexing in the wavenumber domain was proposed to support general communication scenarios.
@@
By ultilizing the EM focusing capability of H-MIMO, efficient near-field WDT and WET were achieved, respectively. Specifically, zhang et. al. [17] investigated a near-field WDT system. The near-field focusing is achieved by fully-digital MIMO, hybrid digital-analog MIMO and H-MIMO, while the H-MIMO performs the best among three structures with a given aperture in the single-user situation. Besides, Wei et. al. [24] considered a tri-polarized H-MIMO aided near-field WDT system. Based on the constructed channel model, two beamforming schemes were proposed to mitigate the interference and improve spectral efficiency. Wu et. al. [25] investigated a H-MIMO assisted multi-user WPT system, where the phase and amplitude of H-MIMO were jointly optimized to generate multifocal beams, in order to improve the transmission efficiency. Zhang et. al. [26] proposed a design of dual-polarized reflective H-MIMO, where near-field focus was realized in a WPT system. Moreover, Zhao et. al. [27] designed a dual-frequency H-MIMO for IDET systems, where each frequency corresponds to a different topology structure in the same aperture.
@@ I-C Motivation and Contributions
  • $\bullet$
    The Friss formula is usually used to calculate the harvested energy in WPT. However, it fails in the near-field range since the formula is under the far-field approximation. In addition, the effect of the beamforming gain and antenna polarization should be further explored in the energy harvesting model.
  • $\bullet$
@@
  • $\bullet$
    Numercial results evaluates the performance of our H-IDET system, which also demonstrate that, the H-IDET scheme outperforms the benchmarks on both near-field focusing and spatial filtering. Besides, the near-field focusing using EM channel model achieves better performance compared to that using the traditional channel model, especially for WPT where the EUs are usually close to the transmitter.
@@
Notation: $(\cdot)^{H}$ denotes conjugate transpose operations; $\mathbb{E}\{\cdot\}$ represents the statistical expectation; $|a|$ and $||\mathbf{a}||$ are the magnitude and norm of a scalar $a$ and vector $\mathbf{a}$, respectively; $||\mathbf{A}||$ denotes the Frobenius norm of the matrix $\mathbf{A}$; $\mathbf{A}(i,j)$ represents the specific element in the $i$-th row and $j$-th column of $\mathbf{A}$; ‘:=’ is signified as ‘is defined to equal’.
@@ II-A H-IDET Transmitter
The transmitter architecture of H-IDET is portrayed in Fig. 1 , which works in a 3-D infinite and homogeneous medium with aperture ${\mathcal{S}}_{\text{T}}$ of area $A_{\text{T}}=|{\mathcal{S}}_{\text{T}}|$. It is assumed that the H-IDET transmitter has continuous antenna aperture and can synthesize any current distribution on its surface. Without loss of generality, we assume that the system works in the narrow band, while the current density distribution (amplitude and phase) at position $\mathbf{s}:=(s_{x},s_{y},s_{z})\in\mathbb{R}^{3\times 1}$ is denoted as $\mathbf{j(s)}\in\mathbb{C}^{3\times 1},\mathbf{s}\in{\mathcal{S}}_{\text{T}}$. A downlink H-IDET system is studied, where the transmitter serves $K$ data users (DUs) and $L$ energy users (EUs), which are all equipped with holographic receivers simultaneously. Let $\mathcal{D}$ and $\mathcal{E}$ represent the set of DUs and EUs, respectively. Then, the set of all users is given by $\mathcal{U}=\mathcal{D}\cup\mathcal{E}$. The symbols transmitted to the users are denoted as $\mathbf{x}\stackrel{{\scriptstyle\vartriangle}}{{=}}[x_{1},\dots,x_{K},x_{K+1},\dots,x_{K+L}]^{T}\in\mathbb{C}^{K+L}$, which are assumed as Gaussian distributed and have normalized power, i.e., $\mathbb{E}\{\mathbf{x}\mathbf{x}^{H}\}=\mathbf{I}_{K+L}$. Similar to the conventional muti-user beamforming, $\bm{\theta}_{k}(\mathbf{s}),{\mathbf{s}}\in{{\mathcal{S}}_{\text{T}}}$ is denoted as the beamformer that H-IDET uses to transmit the symbol $x_{k}$ to the user $k\in\mathcal{U}$. For the simplicity, we assume that beamformers $\{\bm{\theta}_{k}(\mathbf{s})\}^{K}_{k=1}$ are linearly superposed by the H-IDET transmitter, then the current distribution $\mathbf{j(s)}$ can be modeled as [23]
@@ II-B Electromagnetic Channel
where $\kappa=2\pi/\lambda$ is the spatial wavenumber, $\mathsf{i}$ is the imaginary unit and $Z_{0}$ is the characteristic impedance of spatial medium, which is 376.73 Ω in the free space. With the aid of the Green’s method [28] for solving (2), the electric field $\mathbf{e(r)}$ is given by
@@
${\mathbf{G}}({\mathbf{r}},{\mathbf{s}})=\frac{\mathsf{i}\kappa Z_{0}}{4\pi}\frac{e^{\mathsf{i}\kappa\|{\mathbf{p}}\|}}{\|{\mathbf{p}}\|}\left[\left(\mathbf{I_{3}}-\hat{\mathbf{p}}\hat{\mathbf{p}}^{H}\right)+\frac{\mathsf{i}}{\kappa\|{\mathbf{p}}\|}\left(\mathbf{I_{3}}-3\hat{\mathbf{p}}\hat{\mathbf{p}}^{H}\right)+\frac{\mathsf{1}}{(\kappa\|{\mathbf{p}}\|)^{2}}\left(\mathbf{I_{3}}-3\hat{\mathbf{p}}\hat{\mathbf{p}}^{H}\right)\right].$ (4)
@@ II-D Holographic Receiver for EUs
$S(\mathbf{r})$ $=\frac{1}{2Z}\mathbf{e}^{H}_{l}(\mathbf{r})\mathbf{e}_{l}(\mathbf{r})$
$=\frac{1}{2Z}\left\|\sum^{K+L}_{j=1}x_{j}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{l}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right\|^{2}$
@@
$P_{l}$ $=\int_{\mathcal{S}_{\mathrm{R}}}\mathbf{S}(\mathbf{r})\cdot\hat{\bm{\eta}}\mathrm{d}\mathbf{s}$
$=\frac{A_{\text{R}}\cos(\phi)}{2Z}\sum^{K+L}_{j=1}\left\|\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{l}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right\|^{2},$ (9)
@@
Moreover, the non-linear energy harvesting model is conceived in our work, where the energy harvesting power at the EUs can be modeled as [8]
$\Xi(P)=\dfrac{M}{X(1+\exp(-a(P-b)))}-Y$ (10)

### v2

@@ Abstract
Thanks to the application of metamaterials, holographic multiple-input multiple-output (H-MIMO) is expected to achieve a higher spatial diversity gain by enabling the ability to generate any current distribution on the surface. With the aid of electromagnetic (EM) manipulation capability of H-MIMO, integrated data and energy transfer (IDET) system can fully exploits the EM channel to realize energy focusing and eliminate inter-user interference, which yields the concept of holographic IDET (H-IDET). In this paper, we invetigate the beamforming designs for H-IDET systems, where the sum-rate of data users (DUs) are maximized by guaranteeing the energy harvesting requirements of energy users (EUs). In order to solve the non-convex functional programming, a block coordinate descent (BCD) based scheme is proposed, wherein the Fourier transform and the equivalence between the signal-to-interference-plus-noise ratio (SINR) and the mean-square error (MSE) are also conceived, followed by the successive convex approximation (SCA) and an initialization scheme to enhance robustness. Numerical results illustrate that our proposed H-IDET scheme outperforms benchmark schemes, especially the one adopting traditional discrete antennas. Besides, the near-field focusing using EM channel model achieves better performance compared to that using the traditional channel model, especially for WET where the EUs are usually close to the transmitter.
@@ I-A Background
In the era of next Internet of Things (IoT), ubiquitous intelligent connections are establishing between people and things. Deploying millions or even billions of IoT devices having low hardware size, complexity and power consumption into smart cities will readily breeds some emerging applications such as intelligent transportation, virtual reality and cloud computing [1]. However, it is posing a challenge for the energy requirement of all the IoT devices, where manually replacing or charging batteries requires much high maintenance costs and even brings about some safety hazards. In the Ambient IoT, the low-power devices are expected to harvest energy by light, motion, heat and especially radio frequency (RF) signals [2]. By utilizing existing transmitters such as cellular base stations, TV towers, and Wi-Fi access points, RF based wireless energy transfer (WET) has the advantages of better controllability and stability [3]. By coordinating with wireless data transfer (WDT), integrated data and energy transfer (IDET) is a promising technology in the future to provide both energy self-sustaining and data communication services towards low-power devices in IoT [4].
@@
Moreover, thanks to the ability to manipulate the EM field, H-MIMO can not only ultimately approaching the communication limit but also achieve unprecedented concentration of EM energy, which is beneficial to both the WDT and WET [7]. With the aid of the propagation characteristics of the EM channel, the unprecedented energy concentration and multiplexing are realized for EUs, while the inter-user interference is well eliminated for DUs, which yields the concept of Holographic IDET (H-IDET).
@@ I-B Related works
In the multi-user IDET systems, the interference signals are usually harmful for WDT but can improve the WET performance. Therefore, the transceivers should be carefully designed to facilitate efficient IDET. Specifically, Xiong et al. [8] investigated a broadcasting IDET system and analyzed the tradeoff between energy transfer versus information rate under nonlinear energy harvesting model. Yue et al. [9] studied a millimeter wave IDET multicast system, where the data rate was maximized by satisfying wireless charging requirement with the aid of hybrid transceiver design. Large-scale antenna array are used in IDET to alleviate the inherent power attenuation caused by path loss in the channel. Gao et al. [10] investigated an active intelligent reflecting surface (IRS) aided IDET system, where the beamforming designs were obtained by iterative algorithms, in order to maximize the sum-rate of DUs and sum-power harvested by EUs. Zhang et al. [11] considered an IRS assisted IDET system, where a deep reinforcement learning-based approach is proposed by utilizing proximal policy optimization framework, in order to maximize the energy efficiency.
@@
Moreover, there are lots of works studying H-MIMO. The metamaterial element of H-MIMO should satisfy the frequency response of Lorentzian form, which leads to different hardware structures [12]. Precisely, Deng et al. [13] proposed a hardware design for H-MIMO to achieve amplitude-controlled beamforming, where a prototype was implemented to support real-time video transmission. Based on the amplitude-only-controlled hardware structure, they investigated the H-MIMO for cell communication systems [14], low-earth-orbit satellite systems [15] and integrate sensing and communication (ISAC) systems [5]. Besides, different from amplitude-control, the elements of the phase-only-controlled H-MIMO has Lorentzian-constrained phase model while keeping the amplitude constant, which can achieve considerable performance [16, 17]. In addition, J. An et al. [18] investigated a stacked intelligent metasurfaces (SIM) aided H-MIMO communication system, where the novel SIM can accomplish transmit precoding and receiver combining in the native EM wave regime without excessive RF chains.
@@
Furthermore, in order to explore the performance limits enabled by H-MIMO, EM theories should be relied upon for the physical antenna design, communication capacity analysis and channel modelling [19, 20, 21]. Specifically, Badawe et al. [22] proposed a practical physical architecture for the H-MIMO, where experimental results of the fabricated prototype matche well with numerical simulations. Besides, Hu et al. [23] considered a communication system with a H-MIMO transmitter and several single-antenna receivers, where the normalized capacity and Degree of Freedom (DoF) measured per unit surface or per unit volume was analyzed. Moreover, Dardari et al. [24] investigated the communication between two H-MIMO transceivers and the spatial DoF is derived in analytical expression. Furthermore, Pizzo et al. [25] investigated the channel model of H-MIMO in scattering environments, where the model is further applied to evaluate the capacity of H-MIMO communication. In recent years, the EM based H-MIMO transceiver designs for communication systems are up to date. Specifically, Sanguinetti et al. [26] investigated the line-of-sight communication between two H-MIMO transceivers in the near-field. By utilizing Fourier basis functions, a multiplexing scheme was designed to generate orthogonal beamformers. Besides, Zhang et al. [27] considered a multi-user H-MIMO system, where a pattern-division multiplexing in the wavenumber domain was proposed to support general communication scenarios.
@@
By utilizing the EM focusing capability of H-MIMO, efficient near-field WDT and WET were achieved, respectively. Specifically, zhang et al. [17] investigated a near-field WDT system. The near-field focusing is achieved by fully-digital MIMO, hybrid digital-analog MIMO and H-MIMO, while the H-MIMO performs the best among three structures with a given aperture in the single-user situation. Moreover, Sha et al. [28] investigated a H-MIMO assisted WDT system, where a three-dimensional near-field beamforming design is proposed to significantly reduce codebook-size and can be compatible with the fifth-generation new-radio (5G-NR) system. Besides, Wei et al. [29] considered a tri-polarized H-MIMO aided near-field WDT system. Based on the constructed channel model, two beamforming schemes were proposed to mitigate the interference and improve spectral efficiency. Wu et al. [30] investigated a H-MIMO assisted multi-user WET system, where the phase and amplitude of H-MIMO were jointly optimized to generate multifocal beams, in order to improve the transmission efficiency. Zhang et al. [31] proposed a design of dual-polarized reflective H-MIMO, where near-field focus was realized in a WET system. Moreover, Zhao et al. [32] designed a dual-frequency H-MIMO for IDET systems, where each frequency corresponds to a different topology structure in the same aperture.
@@ I-C Motivation and Contributions
  • The Friss formula is usually used to calculate the harvested energy in WET. However, it fails in the near-field range since the formula is under the far-field approximation. In addition, the effect of the beamforming gain and antenna polarization should be further explored in the energy harvesting model.
@@
  • Numerical results evaluates the performance of our H-IDET system, which also demonstrate that, the H-IDET scheme outperforms the benchmarks on both near-field focusing and spatial filtering. Besides, the near-field focusing using EM channel model achieves better performance compared to that using the traditional channel model, especially for WET where the EUs are usually close to the transmitter.
@@
Notation: $(\cdot)^{H}$ denotes conjugate transpose operations; $\mathbb{E}\{\cdot\}$ represents the statistical expectation; $|a|$ and $||\mathbf{a}||$ are the magnitude and norm of a scalar $a$ and vector $\mathbf{a}$, respectively; $||\mathbf{A}||$ denotes the Frobenius norm of the matrix $\mathbf{A}$; $\mathbf{A}(i,j)$ represents the specific element in the $i$-th row and $j$-th column of $\mathbf{A}$; ‘:=’ is signified as ‘is defined to equal’; $\mathfrak{R}(\cdot)$ denotes the real part of its argument.
@@ II-A H-IDET Transmitter
The transmitter architecture of H-IDET is portrayed in Fig. 1 , which works in a 3-D infinite and homogeneous medium with aperture ${\mathcal{S}}_{\text{T}}$ of area $A_{\text{T}}=|{\mathcal{S}}_{\text{T}}|$. It is assumed that the H-IDET transmitter has continuous antenna aperture and can synthesize any current distribution on its surface. Without loss of generality, we assume that the system works in the narrow band, while the current density distribution (amplitude and phase) at position $\mathbf{s}:=(s_{x},s_{y},s_{z})\in\mathbb{R}^{3\times 1}$ is denoted as $\mathbf{j(s)}\in\mathbb{C}^{3\times 1},\mathbf{s}\in{\mathcal{S}}_{\text{T}}$. A downlink H-IDET system is studied, where the transmitter serves $K$ data users (DUs) and $L$ energy users (EUs), which are all equipped with holographic receivers simultaneously. Let $\mathcal{D}$ and $\mathcal{E}$ represent the set of DUs and EUs, respectively. Then, the set of all users is given by $\mathcal{U}=\mathcal{D}\cup\mathcal{E}$. The symbols transmitted to the users are denoted as $\mathbf{x}\stackrel{{\scriptstyle\vartriangle}}{{=}}[x_{1},\dots,x_{K},x_{K+1},\dots,x_{K+L}]^{T}\in\mathbb{C}^{K+L}$, which are assumed as Gaussian distributed [^1] The channel capacity is able to be achieved with the Gaussian distributed transmit signals, which reveals the upper-bound performance of the practical H-IDET system. Note that the Gaussian distributed signals can be asymptotically generated by appropriately designing the modulation and coding schemes for WDT and WET. Specifically, the uncorrelatedness of different WDT streams can be assured by assigning orthogonal codes to different DUs [33]. Moreover, by using pseudo random codes to generate the WET waveforms, they are assured to be uncorrelated with other WET waveforms or with the WDT streams [34]. and have normalized power, i.e., $\mathbb{E}\{\mathbf{x}\mathbf{x}^{H}\}=\mathbf{I}_{K+L}$. Similar to the conventional muti-user beamforming, $\bm{\theta}_{k}(\mathbf{s}),{\mathbf{s}}\in{{\mathcal{S}}_{\text{T}}}$ is denoted as the beamformer that H-IDET uses to transmit the symbol $x_{k}$ to the user $k\in\mathcal{U}$. For the simplicity, we assume that beamformers $\{\bm{\theta}_{k}(\mathbf{s})\}^{K}_{k=1}$ are linearly superposed by the H-IDET transmitter, then the current distribution $\mathbf{j(s)}$ can be modeled as [27]
@@
Remark 1
The Beamformer for each user is essentially a different current distribution over the whole aperture, which the signal of each user is modulated on. For simplicity, it is assumed that the transmitter employs linear superposition to combine multiple information-carrying current distribution [27]. In practice, there are practical hardware architectures that are consistent with our assumptions. For example, in [35] the signal is processed in baseband, then linear superposed and transmitted to the whole aperture by the RF chain and feeds. Therefore, each user can utilize the whole transmitter aperture by the beamformer, and the transmitter is not pre-divided into multiple groups. However, in order to obtain better H-IDET performance, the beamformer of each DU tends to utilize the aperture at different positions and has different phases. The beamformer of each EU tends to have the same current distribution.
@@ II-B Electromagnetic Channel
where $\kappa=2\pi/\lambda$ is the spatial wavenumber, $\lambda$ is the wavelength, $\mathsf{i}$ is the imaginary unit and $Z_{0}$ is the characteristic impedance of spatial medium, which is 376.73 Ω in the free space. With the aid of the Green’s method [36] for solving (2), the electric field $\mathbf{e(r)}$ is given by
@@
${\mathbf{G}}({\mathbf{r}},{\mathbf{s}})$ $=\frac{\mathsf{i}\kappa Z_{0}}{4\pi}\frac{e^{\mathsf{i}\kappa\|{\mathbf{p}}\|}}{\|{\mathbf{p}}\|}\bigg{[}\left(\mathbf{I_{3}}-\hat{\mathbf{p}}\hat{\mathbf{p}}^{H}\right)$
$+\frac{\mathsf{i}}{\kappa\|{\mathbf{p}}\|}\left(\mathbf{I_{3}}-3\hat{\mathbf{p}}\hat{\mathbf{p}}^{H}\right)+\frac{\mathsf{1}}{(\kappa\|{\mathbf{p}}\|)^{2}}\left(\mathbf{I_{3}}-3\hat{\mathbf{p}}\hat{\mathbf{p}}^{H}\right)\bigg{]}.$ (4)
@@ II-D Holographic Receiver for EUs
$S(\mathbf{r})=\frac{1}{2Z}\mathbf{e}^{H}_{l}(\mathbf{r})\mathbf{e}_{l}(\mathbf{r})$ $=\frac{1}{2Z}\left\|\sum^{K+L}_{j=1}x_{j}\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{l}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right\|^{2}$
@@
$P_{l}$ $=\int_{\mathcal{S}_{\mathrm{R}}}\mathbf{S}(\mathbf{r})\cdot\hat{\bm{\eta}}\mathrm{d}\mathbf{s}=\frac{A_{\text{R}}\cos(\phi)}{2Z}\sum^{K+L}_{j=1}\left\|\int_{\mathcal{S}_{\mathrm{T}}}\mathbf{G}_{l}(\mathbf{s})\bm{\theta}_{j}(\mathbf{s})\mathrm{d}\mathbf{s}\right\|^{2},$ (9)
@@
Moreover, the non-linear energy harvesting model is conceived in our work, where the energy harvesting power at the $l$-th EU can be modeled as [8]
$\Xi(P_{l})=\dfrac{M}{X(1+\exp(-a(P_{l}-b)))}-Y$ (10)